OpenStudy (anonymous):

Simplify the expression. 4 + 2 • 8 A)48 b)40 c)20 d)18

OpenStudy (anonymous):

c)20 proof: 2*8=16+4=20

OpenStudy (plainntall):

What are the rules to decide what operation is done first? Do you multiply first or add?

OpenStudy (anonymous):

yes PEMDAS Parenthesis Exponets Multiply and Divide left to right the Add and Subtract left to right

OpenStudy (plainntall):

So ethan's answer is correct. First you multiply 2*8 to get 16 then add 4 to it to get 20. :)

OpenStudy (anonymous):

ya thats right